OpenStudy (anonymous):

Please let me know if what I did is correct: Write an equation of the line parallel to 2y+ 5x -2 = 0 and passes through (-1,2)

OpenStudy (zale101):

First! Write the original equation in slope-intercept form (y=mx+b).

OpenStudy (zale101):

can u do that?

OpenStudy (anonymous):

My answer is: \[y = -\frac{ 5 }{ 2 } x - 5\]

OpenStudy (anonymous):

I posted the answer to the entire question. Hopefully I'm correct.

OpenStudy (zale101):

how did u got this?

OpenStudy (anonymous):

Im going to post what I did: 1m.

OpenStudy (zale101):

the answer supposed to be y=-5/2-1/2

OpenStudy (anonymous):

What did I do wrong then?

OpenStudy (zale101):

u were supposed to do the point slope formula by putting (x1,y1) or (1,2) into the equation y-y1=m(x-x1).

OpenStudy (zale101):

the slope of parallel lines are the same (m1=m2). So, use-5/2 for your slope (m)

OpenStudy (zale101):